OpenStudy (anonymous):

6b=222 7m=84 150=3t 9s=99

OpenStudy (anonymous):

b=37 m=12 t=50 s=33

OpenStudy (gw2011):

6b=222 b=222/6=37 7m=84 m=84/7=12 3t=150 3t=150/3=50 9s=99 9s=99/9=11